Make capital-budgeting decision using net present value (L.O. 5)

Slow to Change Company has decided to computerize its accounting system. The company has two alternatives—it can lease a computer under a three-year contract or purchase a computer outright.

If the computer is leased, the lease payment will be $5,000 each year. The first lease payment will be due on the day the lease contract is signed. The other two payments will be due at the end of the first and second years. The lessor will provide all repairs and maintenance.

If the company purchases the computer outright, it will incur the following costs:$10,500 Acquisition cost Repairs and maintenance: First year Second year Third year 300 250 350

The computer is expected to have only a three-year useful life because of obsolescence and technological advancements. The computer will have no salvage value and be depreciated on a double-declining-balance basis. Slow to Change Company’s cost of capital is 16%.

a.   Calculate the net present value of out-of-pocket costs for the lease alternative.
b.   Calculate the net present value of out-of-pocket costs for the purchase alternative.
c.   Do you recommend that the company purchase or lease the machine?
Answer #1

(a) NPV of Lease alternative :-

= 5,000 + [5,000 x (PVAF@16%,2 years)]

= 5,000 + (5,000 x 1.6052)

= $13,026

(b) NPV of Purchase alternative :-

= 10,500 + (300 x PVF@16%,1 year) + (250x PVF@16%,2 year) + (350 x PVF@16%,3 year)

= 10,500 + (300 x 0.8621) + (250 x 0.7432) + (350 x 0.6407)

= $11,169

(c) It is recommended to Purchase the machine instead of leasing due to lesser out of pocket cost in purchase alternative.